# Affective Computing in the Era of Large Language Models: A Survey from the NLP Perspective

### v1

@@ Abstract
Affective Computing (AC), integrating computer science, psychology, and cognitive science knowledge, aims to enable machines to recognize, interpret, and simulate human emotions. To create more value, AC can be applied to diverse scenarios, including social media, finance, healthcare, education, etc. Affective Computing (AC) includes two mainstream tasks, i.e., Affective Understanding (AU) and Affective Generation (AG). Fine-tuning Pre-trained Language Models (PLMs) for AU tasks has succeeded considerably. However, these models lack generalization ability, requiring specialized models for specific tasks. Additionally, traditional PLMs face challenges in AG, particularly in generating diverse and emotionally rich responses. The emergence of Large Language Models (LLMs), such as the ChatGPT series and LLaMA models, brings new opportunities and challenges, catalyzing a paradigm shift in AC. LLMs possess capabilities of in-context learning, common sense reasoning, and advanced sequence generation, which present unprecedented opportunities for AU. To provide a comprehensive overview of AC in the LLMs era from an NLP perspective, we summarize the development of LLMs research in this field, aiming to offer new insights. Specifically, we first summarize the traditional tasks related to AC and introduce the preliminary study based on LLMs. Subsequently, we outline the relevant techniques of popular LLMs to improve AC tasks, including Instruction Tuning and Prompt Engineering. For Instruction Tuning, we discuss full parameter fine-tuning and parameter-efficient methods such as LoRA, P-Tuning, and Prompt Tuning. In Prompt Engineering, we examine Zero-shot, Few-shot, Chain of Thought (CoT), and Agent-based methods for AU and AG. To clearly understand the performance of LLMs on different Affective Computing tasks, we further summarize the existing benchmarks and evaluation methods. Finally, we discuss challenges from an NLP perspective, such as ethical considerations and effective evaluation, and outline future directions. Our survey seeks to provide valuable insights and references for researchers in this rapidly evolving domain.
Index Terms:
Affective Computing, Large Language Models, Instruction Tuning, Prompt Engineering, Benchmarks.
@@ 1 Introduction
Affective Computing (AC), encompassing the study and development of systems capable of recognizing, interpreting, processing, and simulating human affects through natural language and other modalities, seeks to bridge the gap between human emotions and machine understanding [1, 2, 3]. The AC task can be broadly categorized into two mainstream tasks: the Affective Understanding (AU) task [4] and the Affective Generation (AG) task [5, 6]. Affective Understanding focuses on recognizing and interpreting human emotions, covering Sentiment Analysis [7, 8], Sarcasm Detection [9, 10], and so on. Affective Generation involves creating responses or content that can elicit specific emotional responses in humans to generate expressions with emotionally nuanced, including Emotional-aware Dialogue Generation[11, 12], Creative Content Generation, etc. With the emergence of Large Language Models (LLMs)[13, 14, 15], the Affective Computing task is experiencing a paradigm shift to explore and model human emotions in unprecedented ways.
Figure 1: Comparison of approaches to AC Tasks in the PLMs era and the LLMs era. In the PLMs era, specific datasets were typically used to fine-tune PLMs. Each specific dataset usually requires separate model fine-tuning to achieve optimal performance. When addressing AC tasks, PLMs often utilized sentiment-aware modules to assist in AU and AG tasks. The deep neural networks (DNN) [16] used for AC tasks could have various structures. For example, a multilayer perceptron (MLP) [17] might be used for AU tasks, while a long short-term memory network (LSTM) [18] might be used for AG tasks. In contrast, LLMs possess strong generalization capabilities, allowing them to handle multiple tasks simultaneously without needing for separate fine-tuning for each specific task. When using adapters for instruction tuning, the parameters of the LLM are typically frozen to reduce resource consumption and improve efficiency.
@@
Prior to the advent of LLMs, the dominant paradigm of Affective Computing involves fine-tuning Pre-trained Language Models (PLMs) [19], such as BERT [20] and RoBERTa [21], on labelled datasets for specific tasks to improve performance. While the above training paradigm has achieved numerous successes in the tasks related to Affective Understanding, it still is inherently limited by the quality and scale of manually annotated data and struggles to generalize to new domains [22]. The rise of LLMs research, as demonstrated in Figure 1, profoundly transforms the research paradigm of AC, expanding task variety and surpassing the constraints of previous studies, which often focus on particular datasets and single-task the performance improvements [23, 24]. This evolution is accompanied by developing novel approaches that exploit the vast knowledge of LLMs [25] to simultaneously boost the performance of various tasks related to Affective Computing.
@@
The auto-regressive nature and massive pre-training of large language models give them an advantage in processing dialogue inputs and generating coherent responses [26, 27]. Various techniques are proposed to fully harness and enhance the capabilities of LLMs, including instruction tuning and prompt engineering. Specifically, Instruction Tuning [28, 29] involves fine-tuning a pre-trained language model on a dataset where the tasks are explicitly defined through instructions. For example, Flan-T5 [28] specific designs instructions for different tasks, e.g., Sentiment Analysis, Commonsense Reasoning, Question Answering, and so on. Instruction Tuning is a form of targeted training that aims to make the model more adept at following complex and varied instructions across different domains. Prompt Engineering [30, 31], on the other hand, is more about crafting the input given to the model to elicit the desired output without modifying the model itself. In-context Learning can further mine a model to learn and apply knowledge based on the instances provided directly within the prompt, without any prior explicit training on these specific tasks [32]. Although LLMs offer new opportunities and show initial success in various tasks, a comprehensive review that consolidates and summarizes the recent advancements in AC with LLMs from an NLP perspective still needs to be present. Existing surveys either focus narrowly on specific tasks like sentiment analysis or a limited set of classic affective computing tasks [23, 33, 34, 35, 36], or they concentrate on traditional, non-LLM-based methods [19, 37].
@@
To mitigate the gap, it is essential to provide a comprehensive overview of how LLMs are being applied across the diverse spectrum of tasks related to AC, along with the unique opportunities and challenges they bring. Accordingly, we investigate the capabilities and limitations of LLMs in AC tasks, covering both Affective Understanding and Generation. While our focus is primarily on NLP, we also include text-centric multimodal affective computing tasks, recognizing the growing importance of integrating multiple modalities in AC. We begin by summarizing early studies that laid the groundwork for understanding the capabilities and limitations of LLMs in AC. Building on this, we explore two critical technologies for adapting LLMs to AC tasks: instruction tuning and prompt engineering. It involves reviewing various fine-tuning methods designed to enhance LLMs’ performance on AC tasks, followed by an exploration of different prompting strategies aimed at eliciting desired emotional responses from LLMs. To deliver a comprehensive assessment, we further summarize benchmarks and evaluation methods for systematically evaluating LLMs’ performance in AC, including the innovative use of LLMs as evaluation tools. Finally, we enumerate open challenges and outline promising future research directions, offering insights into this rapidly advancing field’s continuous evolution and potential.
@@ 2 Tasks
The mainstream tasks of Affective Computing (AC) include Affective Understanding (AU) [4] and Affective Generation (AG) [5, 6]. The primary traditional tasks and related subtasks of AU and AG are summarized in Table I.
TABLE I: The traditional tasks of Affective Computing from an NLP perspective, with representative methods in the LLM era.
AC Tasks Subtasks Detailss
AU Sentiment Analysis (SA) Polarity Classification (PC) USA [38], WisdoM [39], RP-CoT [40], Sun et.al [41]
Emotional Classification (EC) Zhang et.al [35], Stigall et.al [42]
Aspected-Based Sentiment Analysis (ABSA) InstructABSA [43], SCRAP [44]
Emotional Intensity Detection (EID) Amin et.al [45]
Implicit Sentiment Analysis (ISA) THOR [46]
Emotion Recognition in Conversation (ERC) UniMSE [47], DialogueLLMzhang2023dialoguellm, CKERC [48], InstructERC [49]
Emotion Cause Paired Extract (ECPE) DECC [50], ECR-Chain [51]
Subjective Text Analysis (STA) Suicide Tendency Detection (STD) Zhang et.al [35], Amin et.al [45]
Personality Assessment (PA) Zhang et.al [35], Amin et.al [45]
Toxicity Detection (TD) Zhang et.al [35], Amin et.al [45]
Sarcasm Detection (SD) Hoffmann et.al [52], Zhang et.al [35], Amin et.al [45]
Well-Being Assessment (WBA) Zhang et.al [35], Amin et.al [45]
Engagement Measurement (EM) Zhang et.al [35], Amin et.al [45]
AG Emotional Dialogue (ED) Empathetic Response Generation (ERG) Sibyl [53], PromptMind [54], PEGS [55], Chain of Empathy [56], ECoT [57]
Emotional Support Conversation (ESC) SoulChat [58], EmoLLM [59], PICA [60], Cue-CoT [61]
Review Summarization (RS) Opinion Summarization (OS) TCG [62], LASS [63]

### v2

@@ Abstract
Affective Computing (AC) integrates computer science, psychology, and cognitive science to enable machines to recognize, interpret, and simulate human emotions across domains such as social media, finance, healthcare, and education. AC commonly centers on two task families: Affective Understanding (AU) and Affective Generation (AG). While fine-tuned pre-trained language models (PLMs) have achieved solid AU performance, they often generalize poorly across tasks and remain limited for AG, especially in producing diverse, emotionally appropriate responses. The advent of Large Language Models (LLMs) (e.g., ChatGPT and LLaMA) has catalyzed a paradigm shift by offering in-context learning, broader world knowledge, and stronger sequence generation. This survey presents an NLP-oriented overview of AC in the LLM era. We (i) consolidate traditional AC tasks and preliminary LLM-based studies; (ii) review adaptation techniques that improve AU/AG, including Instruction Tuning (full and parameter-efficient methods such as LoRA, P-/Prompt-Tuning), Prompt Engineering (zero/few-shot, chain-of-thought, agent-based prompting), and Reinforcement Learning. For the latter, we summarize RL from human preferences (RLHF), verifiable/programmatic rewards (RLVR), and AI feedback (RLAIF), which provide preference- or rule-grounded optimization signals that can help steer AU/AG toward empathy, safety, and planning, achieving finer-grained or multi-objective control. To assess progress, we compile benchmarks and evaluation practices for both AU and AG. We also discuss open challenges—from ethics, data quality, and safety to robust evaluation and resource efficiency—and outline research directions. We hope this survey clarifies the landscape and offers practical guidance for building affect-aware, reliable, and responsible LLM systems.
keywords:
[…]
@@ 1 Introduction
Affective Computing (AC), encompassing the study and development of systems capable of recognizing, interpreting, processing, and simulating human affects through natural language and other modalities, seeks to bridge the gap between human emotions and machine understanding Picard (2000); Tao & Tan (2005); Banafa (2024). The AC task can be broadly categorized into two mainstream tasks: the Affective Understanding (AU) task Zhao et al. (2023a) and the Affective Generation (AG) task Nie & Zhan (2022); Pei et al. (2024). Affective Understanding focuses on recognizing and interpreting human emotions, covering Sentiment Analysis Medhat et al. (2014); Taboada (2016), Sarcasm Detection Joshi et al. (2017); Verma et al. (2021), and so on. Affective Generation involves creating responses or content that can elicit specific emotional responses in humans to generate expressions with emotionally nuanced, including Emotional-aware Dialogue Generation Rashkin et al. (2019); Liu et al. (2021), Creative Content Generation, etc. With the emergence of Large Language Models (LLMs) OpenAI (2023); Touvron et al. (2023); Zeng et al. (2023), the Affective Computing task is experiencing a paradigm shift to explore and model human emotions in unprecedented ways.
Figure 1: Comparison of approaches to AC tasks across the PLMs and LLMs eras. In the PLMs era, task-specific datasets fine-tune PLMs with sentiment-aware modules and downstream DNNs Bengio et al. (2009) (e.g., MLP Rosenblatt (1958), LSTM Hochreiter & Schmidhuber (1997)), typically requiring a separate model per task. In the LLMs era, a single (mostly frozen) LLM handles multiple tasks via (i) instruction tuning with adapters, (ii) prompt engineering, and (iii) reinforcement learning that samples rollouts and uses reward/preference signals to optimize a loss and update the adapters.
@@
Prior to the advent of LLMs, the dominant paradigm of Affective Computing involves fine-tuning Pre-trained Language Models (PLMs) Amin et al. (2023a), such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), on labelled datasets for specific tasks to improve performance. While the above training paradigm has achieved numerous successes in the tasks related to Affective Understanding, it still is inherently limited by the quality and scale of manually annotated data and struggles to generalize to new domains Mao et al. (2023). The rise of LLMs research, as demonstrated in Figure 1, profoundly transforms the research paradigm of AC, expanding task variety and surpassing the constraints of previous studies, which often focus on particular datasets and single-task the performance improvements Tan et al. (2023a); Krugmann & Hartmann (2024). This evolution is accompanied by developing novel approaches that exploit the vast knowledge of LLMs Wei et al. (2022a) to simultaneously boost the performance of various tasks related to Affective Computing.
@@
The auto-regressive nature and massive pre-training of LLMs give them an advantage in processing dialogue inputs and generating coherent responses Radford et al. (2018); Brown et al. (2020). Various techniques are proposed to fully harness and enhance the capabilities of LLMs, including Instruction Tuning, Prompt Engineering and Reinforcement Learning. Specifically, Instruction Tuning Chung et al. (2024); Zhang et al. (2024d) involves fine-tuning a pre-trained language model on the dataset where tasks are explicitly defined through instructions. For example, Flan-T5 Chung et al. (2024) designs specific instructions for different tasks, e.g., Sentiment Analysis, Commonsense Reasoning, Question Answering, etc. Instruction Tuning is a form of targeted training that aims to make the model more adept at following complex and varied instructions across different domains. Prompt Engineering Zhou et al. (2023b); White et al. (2023) is more about crafting the input given to the model to elicit the desired output without modifying the model itself. In-context Learning can further mine a model to learn and apply knowledge based on the instances provided directly within the prompt, without any prior explicit training on these specific tasks Xie et al. (2022). Reinforcement Learning optimizes the model against preference-based or verifiable rewards to improve helpfulness, safety, and style. A common pipeline, Reinforcement Learning from Human Feedback (RLHF), first trains a reward model from human comparisons and then updates the policy with online RL—typically Proximal Policy Optimization (PPO) Lambert (2025). Complementary directions include Reinforcement Learning with Verifiable Rewards (RLVR) Lambert et al. (2024), which uses rule-checkers or pass/fail signals as rewards, and methods that reduce or avoid online RL, such as Direct Preference Optimization (DPO) Rafailov et al. (2023a). Reinforcement Learning from AI Feedback (RLAIF) distills feedback from capable AI evaluators or rule sets into the preference signal (e.g., Wen et al. (2025)). For AC tasks, these approaches allow alignment directly toward affect-aware objectives, covering politeness, empathy, and non-toxic style.
@@
Although LLMs offer new opportunities and show initial success in various tasks, a comprehensive review that consolidates and summarizes the recent advancements in AC with LLMs from an NLP perspective still needs to be present. Existing surveys either focus narrowly on specific tasks like sentiment analysis or a limited set of classic affective computing tasks Tan et al. (2023a); Wake et al. (2023); Cui et al. (2023); Zhang et al. (2023e); Zhu et al. (2024), or they concentrate on traditional, non-LLM-based methods Amin et al. (2023a); Afzal et al. (2023). To mitigate the gap, it is essential to provide a comprehensive overview (Figure 2) of how LLMs are being applied across the diverse spectrum of tasks related to AC, along with the unique opportunities and challenges they bring. Accordingly, we investigate the capabilities and limitations of LLMs in AC tasks, covering both Affective Understanding and Generation. While our focus is primarily on NLP, we also include text-centric multimodal affective computing tasks, recognizing the growing importance of integrating multiple modalities in AC. We begin by summarizing early studies that laid the groundwork for understanding the capabilities and limitations of LLMs in AC. Building on this, we explore three critical technologies for adapting LLMs to AC tasks: instruction tuning (§4), prompt engineering (§5) and reinforcement learning (§6). While tuning grounds the model in task formats and prompting controls behavior at inference, RL supplies preference- or programmatically verifiable objectives that close the optimization loop. To deliver a comprehensive assessment, we further summarize benchmarks and evaluation methods (§7) for systematically evaluating LLMs’ performance in AC, including the innovative use of LLMs as evaluation tools. Finally, we enumerate open challenges and outline promising future research directions (§8), offering insights into this rapidly advancing field’s continuous evolution and potential.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=4cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, l sep=12pt, , [Affective Computing, root, [Preliminary Study (§3), preliminary [Affective Understanding (§3.1), preliminary [Tan et al. Tan et al. (2023b); Rodriguez et al. Rodr´ıguez-Ibánez et al. (2023); Das et al. Das & Singh (2023); Krugmann et al. Krugmann & Hartmann (2024); Zhang et al. Zhang et al. (2023f); Belal et al. Belal et al. (2023); Lossio et al. Lossio-Ventura et al. (2024); Alexander et al. Alexandersen & Rutlin (2024); Amin et al. Amin et al. (2023b); Wang et al. Wang et al. (2024h); Wu et al. Wu et al. (2024a); Zhang et al. Zhang et al. (2023d); M.M.Amin et al. Amin et al. (2023a), preliminary, text width=12cm ] ] [Affective Generation (§3.2), preliminary [ Zhao et al. Zhao et al. (2023a); Welivita et al. Welivita & Pu (2024); Elyoseph et al. Elyoseph et al. (2024); SECEU Wang et al. (2023e); EQ-Bench Paech (2024); SOUL Deng et al. (2023d), preliminary, text width=12cm ] ] ] [Instruction Tuning (§4), sft [Parameter-Efficient Fine-Tuning, sft [ EMO-LLaMA Xing et al. (2024); Emotion-LLaMA Cheng et al. (2024b); InstructERC Lei et al. (2024); CKERC Fu (2024); DialogueLLM Zhang et al. (2023h); Iain J. et. al Cruickshank & Ng (2024); STOEI Jia et al. (2024); EKTC Cao et al. (2025); Wang et al. Wang et al. (2023b); PEGS Zhang et al. (2024f); GSA Hou et al. (2024); Penf et al. Peng et al. (2024); Ding et al. Ding et al. (2024); M2SE Li et al. (2024a); MoEI Zhao et al. (2024b); Zhao et al. Zhao et al. (2024b); PICA Zhang et al. (2023i); Zheng et al. Zheng et al. (2023b); BLSP-Emo Wang et al. (2024a); EMOLLM Liu et al. (2024c); FAME-Net Wang et al. (2024e); PerceptiveAgent Yan et al. (2024), sft, text width=12cm] ] [Full-Parameter Fine-Tuning, sft [Colin Raffel et al. Raffel et al. (2020a); Boyu Zhang et al. Zhang et al. (2023a); USA Gan et al. (2023); Wisdom Wang et al. (2024f); Omni-Emotion Yang et al. (2025); InstructABSA Scaria et al. (2023); ITSCL Zhang et al. (2024b); Siddharth Varia et al. Varia et al. (2023); UniSA Li et al. (2023b); NUS-Emo Luo et al. (2024); ObG Wang et al. (2024b); SoulChat Chen et al. (2023b); EMOVA Chen et al. (2024a); Emollms Liu et al. (2024c); MODA Zhang et al. (2025c); Emoada Dong et al. (2024); Lim Dongjun et al. Lim & Cheong (2024), sft, text width=12cm] ] ] [Prompt Engineering (§5), prompt [Zero-shot (§5.1.1 & 5.2.1), prompt [ Amin et al. Amin et al. (2023b); Ouyang et al. Ouyang et al. (2024); M.A.Al Asad et al. Al Asad et al. (2023); W.Stigall et al. Stigall et al. (2024); Hong et al. Hong et al. (2025); Zhou et al. Zhou et al. (2023a); Bai et al. Bai et al. (2024); Cabello et al. Cabello & Akujuobi (2024); Wang et al. Wang & Luo (2023) Niu et al. Niu et al. (2024); Plaza et al. Plaza-del Arco et al. (2024); Li et al. Li et al. (2023a); Shen et al. Shen et al. (2024); Wu et al. Wu et al. (2024d); Santoso et al. Santoso et al. (2024); Liang et al. Liang et al. (2024); Wang et al. Wang et al. (2024g) Lee et al. Lee et al. (2023a); Yang et al. Yang et al. (2024c); TCG Bhaskar et al. (2022); Y.M.Resendiz Resendiz & Klinger (2023); PromptMind Su et al. (2023), prompt, text width=12cm]] [Few-shot (§5.1.2 & 5.2.2), prompt [ Zhou et al. Zhou et al. (2024); Zhang et al. Zhang et al. (2024d); Yang et al. Yang et al. (2024a); SCRAP Kim et al. (2024); Wisdom Yang et al. (2024a); LVLMS Liu et al. (2024a); Ye et al. (2024); SEGA Chen et al. (2024c); Sibyl Wang et al. (2023c); Chen et al. Chen et al. (2023a); Kang et al. Kang et al. (2024); LLaMA2 Touvron et al. (2023), prompt, text width=12cm]] [Chain-of-Thought (§5.1.3 & 5.2.3), prompt [ Lee et al. Lee et al. (2024); THOR Fei et al. (2023); Lyu et al. Lyu et al. (2024); CoE Lee et al. (2023b); Manzoor et al. Manzoor et al. (2024); ECR-Chain Huang et al. (2024); RP-CoT Wang & Luo (2023); DECC Wu et al. (2024b); EMO Baumann & Kramer (2024); Deng Deng et al. (2023a); CFEG Chen et al. (2024b); Cue-CoT Wang et al. (2023a); EDIT Wu et al. (2023); Aptness Hu et al. (2024); ProCoT Deng et al. (2023b); ECoT Li et al. (2024g); ESCoT Zhang et al. (2024e); Li et al. Li et al. (2024d); VEHICLE Shao et al. (2024a), prompt, text width=12cm]] [Agent-based (§5.1.4 & 5.2.4), prompt [ Fei et al. Fei et al. (2023); Yang et al. Yang et al. (2024a); Sun et al. Sun et al. (2023); Wei et al. Wei et al. (2024); E.S.Van et al. Van Haeringen et al. (2023); PANAS Regan et al. (2024); HAD Xing (2024); Agent4SC Zhang et al. (2024f); COOPER Cheng et al. (2024a); PerceptiveAgent Yan et al. (2024); Emotion-aware agents Hu et al. (2022b); Zhang et al. Zhang et al. (2024c), prompt, text width=12cm]] ] [Reinforcement Learning (§6), rl [RLHF(§6.1), rl [ Zhu et al. Zhu et al. (2023a); XU et al. Xu et al. (2025); AlignCap Liang et al. (2024); Chen et al. Chen et al. (2025), rl, text width=12cm] ] [RLVR(§6.2), rl [ R1-Omni Zhao et al. (2025a); SAGE (State Augmented GEneration) Zhang & Jaitly (2025); SAGE (Sentient Agent as a Judge) Zhang et al. (2025a); CRP Li et al. (2024c); EmpRL Ma et al. (2025); ReflectDiffu Yuan et al. (2024); ABLE Mishra et al. (2024); GENTEEL-NEGOTIATOR Priya et al. (2025); F2RL Wang et al. (2024d); Unnikrishnan et al. Unnikrishnan (2024), rl,text width=12cm] ] [RLAIF(§6.3), rl [ Yoshida et al. Yoshida et al. (2025) PPDPP Deng et al. (2023c); DialogXpert Rakib et al. (2025); CSO Zhao et al. (2025b); Ye et al. Ye et al. (2025); Decoupled ESC Zhang et al. (2025b), rl,text width=12cm] ] ] [Benchmark & Evaluation (§7), benchmark [Affective Understanding (§7.1), benchmark [SOUL Deng et al. (2023d); GPT-4v with Emotion Lian et al. (2023b); MERBench Lian et al. (2024); DFEW Jiang et al. (2020); MC-EIU Liu et al. (2024b), benchmark, text width=12cm] ] [Affective Generation (§7.2), benchmark [ Cue-CoT Wang et al. (2023a); ESC-Eval Zhao et al. (2024a); MEDIC Zhu et al. (2023b); EmotionBench tse Huang et al. (2023); STICKERCONV Zhang et al. (2024f); SECEU Wang et al. (2023d); EQ-Bench Paech (2024); EmoBench Sabour et al. (2024); EIBENCH Zhao et al. (2024b); EmoVIT Xie et al. (2024); EMER Lian et al. (2023a); MER-UniBench Lian et al. (2025), benchmark, text width=12cm] ] [Others (§7.3), benchmark [ ChatGPT2AC Amin et al. (2024); MM-INSTRUCTEVAL Yang et al. (2024b), benchmark, text width=12cm] ] ] ]
Figure 2: The overview of the paper.
@@ 2 Tasks
Affective Computing (AC) comprises two mainstream task families: Affective Understanding (AU) Zhao et al. (2023a) and Affective Generation (AG) Nie & Zhan (2022); Pei et al. (2024), as summarized in Table 1.
Table 1: The traditional tasks of Affective Computing from an NLP perspective, with representative methods in the LLM era.
AC Tasks Subtasks Examples
AU Sentiment Analysis (SA) Polarity Classification (PC) USA Gan et al. (2023), WisdoM Wang et al. (2024f), RP-CoT Wang & Luo (2023), Sun et.al Sun et al. (2023)
Emotional Classification (EC) Zhang et.al Zhang et al. (2023e), Stigall et.al Stigall et al. (2024)
Aspect-Based Sentiment Analysis (ABSA) InstructABSA Scaria et al. (2023), SCRAP Kim et al. (2024)
Emotional Intensity Detection (EID) Amin et.al Amin et al. (2023b)
Implicit Sentiment Analysis (ISA) THOR Fei et al. (2023)
Emotion Recognition in Conversation (ERC) UniMSE Hu et al. (2022a), DialogueLLM Zhang et al. (2023h), CKERC Fu (2024), InstructERC Lei et al. (2024)
Emotion Cause Paired Extraction (ECPE) DECC Wu et al. (2024b), ECR-Chain Huang et al. (2024)
Subjective Text Analysis (STA) Suicide Tendency Detection (STD) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Personality Assessment (PA) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Toxicity Detection (TD) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Sarcasm Detection (SD) Hoffmann et.al Hoffmann et al. (2022), Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Well-Being Assessment (WBA) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Engagement Measurement (EM) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
AG Emotional Dialogue (ED) Empathetic Response Generation (ERG) Sibyl Wang et al. (2023b), PromptMind Yang et al. (2024c), PEGS Zhang et al. (2024f), Chain of Empathy Lee et al. (2023b), ECoT Li et al. (2024f)
Emotional Support Conversation (ESC) SoulChat Chen et al. (2023b), EmoLLM EmoLLM (2024), PICA Zhang et al. (2023i), Cue-CoT Wang et al. (2023a)
Review Summarization (RS) Opinion Summarization (OS) TCG Bhaskar et al. (2022), LASS Li et al. (2023a)